# Elicitability and identifiability of tail risk measures

### v2

@@ 4 Identifiability results
To illustrate Theorem 4.3, we let $\rho^{*}$ be the expectation with a strict $\mathcal{M}^{1}$-identification function $V^{*}(x,y)=x-y$, $x,y\in\mathbb{R}$. The corresponding $p$-tail risk measure $\rho$ induced by $\rho^{*}$ is $\mathrm{ES}_{p}$. Indeed, a straightforward calculation yields that the second component of (4.3) is $(1-p)$ times the second component of (3.3).
@@ 5.1 Elicitability relations and score functions
[…]
@@ 5.2 Special cases and discussions
Take the $p$-tail pair $(\rho,\rho^{*})=(\mathrm{ES}_{p},\mathbb{E})$. If we choose the strictly $\mathcal{M}^{1}$-consistent score $S^{*}(x,y)=\frac{1}{1-p}\big{(}-\phi(x)+\phi^{\prime}(x)(x-y)\big{)}+g(y)$ with $\phi$ strictly convex, then the score in (5.4) coincides with the one in (3.4). The condition that $y\mapsto S^{*}(x,y)$ is strictly increasing is equivalent to the condition that $y\mapsto g(y)-\frac{1}{1-p}\phi^{\prime}(x)y$ is strictly increasing. Invoking Lemma 5.5, a necessary and sufficient condition for the existence of such a $g$ is that $\phi^{\prime}$ is bounded from above.
@@
where $\mathcal{P}(y)$ denotes the set of all partitions $P=\{z_{0},z_{1},\ldots,z_{n_{P}}:z_{i}<z_{i+1}\}$ of the interval $[\min(0,y),\max(0,y)]$, then we obtain a strictly consistent score $S$ in (5.4) upon choosing $g(y)=\|u\|(y)+\|t\|(y)+y$, $y\in\mathbb{R}$. For example, if $u=1$ and $t(y)=y^{2}$, then $\|u\|(y)=0$ and $\|t\|(y)=\mathrm{sign}(y)y^{2}$ for $y\in\mathbb{R}$. If $u(y)=\mathds{1}_{[a,b)}(y)$ for $a<b$, then $\|u\|(y)=\mathds{1}_{[a,\infty)}(y)+\mathds{1}_{[b,\infty)}(y)$ for $y\in\mathbb{R}$.
@@ 6 Results on the left tail or the body of the distribution
and a risk measure $\rho^{[p,q]}$ via another risk measure $\rho^{*}$ by
@@
Let $\rho^{*}\colon\mathcal{M}\to\mathbb{R}$ be $\mathcal{M}$-elicitable with strictly $\mathcal{M}$-consistent score $S^{*}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$. For $10<p<q<$ 1, let $\rho^{[p,q]}$ be defined via (6.1). Then the function given by
$S(v_{1},v_{2},x,y)$
@@
Theorem 6.2 generalizes the elicitability result for the Range Value-at-Risk of Fissler and Ziegel (2021b). Identifiability results in the spirit of Theorem 4.3 can also be obtained.
@@ 7 Concluding remarks
Second, we provided in Theorem 5.3 some sufficient conditions on the elicitable risk measure $\rho^{*}$ that guarantee the elicitability of $p$-tail risk measure $\rho$ itself or the pair $(Q_{p},\rho)$. A full characterization of all elicitable risk measures $\rho^{*}$ yielding elicitability of $\rho$ or $(Q_{p},\rho)$ is missing. Theorem 5.3 does not provide a full answer to this question. In particular, one needs to rely on the knowledge of the existence of a strictly consistent scoring function for $\rho^{*}$, $S^{*}(x,y)$, that is strictly increasing in $y$. Third, the elicitation complexity (Frongillo and Kash, 2021) of tail risk measures remains unclear. Roughly (with suitable regularizing conditions), the elicitation complexity of $\rho$ is a positive integer $k$ that quantifies how many dimensions are needed to make $\rho$ a function of an elicitable vector $(\rho_{1},\dots,\rho_{k})$. It may be tempting to conjecture that if $\rho^{*}$ has elicitation complexity $k$, then $\rho$ has elicitation complexity at most $k+1$, seeing from the example of $(\rho,\rho^{*})=(\mathrm{ES}_{p},\mathbb{E})$ with elicitation complexity $(2,1)$; see Frongillo and Kash (2021). This question seems to be very challenging to answer with current techniques.
@@ Acknowledgements
F. Liu and R. Wang acknowledge financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2024-03728, CRC-2022-00141, RGPIN-2020-04717, DGECR-2020-00340). L. Wei acknowledges financial support from the National Natural Science Foundation of China (No. 12271415).
@@ Appendix A Risk measures
Some desirable properties of risk measures are listed below.
@@
    Law-invariance: if $X,Y\in\mathcal{X}$ and $X\buildrel\,\mathrm{d}\over{=}Y$, then $\hat{\rho}(X)=\hat{\rho}(Y)$.
@@
    Comonotonic additivity: $\hat{\rho}(X+Y)=\hat{\rho}(X)+\hat{\rho}(Y)$ if $X,Y\in\mathcal{X}$ are comonotonic.333 Two random variables $X$ and $Y$ are comonotonic if there exists $\Omega_{0}\in\mathcal{F}$ with $\mathbb{P}(\Omega_{0})=1$ and $(X(\omega)-X(\omega^{\prime}))(Y(\omega)-Y(\omega^{\prime}))\geqslant 0~{}~{}\mbox{for all}~{}\omega,\omega^{\prime}\in\Omega_{0}.$ Comonotonicity of $X$ and $Y$ is equivalent to the existence of a random variable $Z\in L^{0}$ and two non-decreasing functions $f$ and $g$ such that $X=f(Z)$ and $Y=g(Z)$ almost surely. We refer to Dhaene et al. (2002) for an overview on comonotonicity.
@@
A risk measure $\rho$ is a monetary risk measure if it satisfies (A1) and (A2), it is a convex risk measure if it satisfies (A1)–(A3), and it is a coherent risk measure if it satisfies (A1)–(A4). The acceptance set of a monetary risk measure $\rho$ is defined as $\mathcal{A}_{\rho}=\{X\in\mathcal{X}:\hat{\rho}(X)\leqslant 0\}.$
@@ Appendix B Assumptions
    The identification function $V\colon\mathbb{R}^{3}\to\mathbb{R}^{2}$ defined by
    $V(v,x,y)=\begin{pmatrix}\mathds{1}_{\{y\leqslant v\}}-p\\ \mathds{1}_{\{y>v\}}V_{\rho^{*}}(x,y)\end{pmatrix}$
    is such that its expectation with respect to any $F\in\mathcal{M}$ is continuously differentiable. In particular, any $F\in\mathcal{M}$ has a continuous derivative $f$ (which is also its Lebesgue density).
@@
    Suppose that the complement of the set
    $\{(v,x,y)\in\mathsf{A}\times\mathbb{R}:V(v,x,\cdot)\text{ and }S(v,x,\cdot)\text{ are continuous at the point }y\}$
    has 3-dimensional Lebesgue measure zero.
@@ Appendix C Proofs in Section 5
    Let $S(v,x,y)$, $v,x,y\in\mathbb{R}$, be a strictly $\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$-consistent score and $S_{r}^{*}$ given in (5.2) for some $r\in\mathbb{R}$. For any $F\in\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$ and $v\in Q_{p}(F)$ it holds that
    $\rho(F)=\operatorname*{arg\,min}_{x\in\mathbb{R}}\int_{\mathbb{R}}S(v,x,y)\,\mathrm{d}F(y).$
    Let $G\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}$. Then, as in part (i), define $F(y)=\mathds{1}_{\{y\geqslant r\}}\big{(}(1-p)G(y)+p\big{)}$, resulting in $F\in\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$ and $r=\mathrm{VaR}_{p}^{-}(F)\in Q_{p}(F)$. Again, $F_{p}=G$, which implies
@@
is a strict $\mathcal{M}$-identification function for $(\mathrm{VaR}_{p}^{-},\rho)$; see also (Fissler and Hoga, 2023, Proposition 3.4). We obtain for $F\in\mathcal{M}$ with derivative (and density) $f$

### v3

@@ 4 Identifiability results
To illustrate Theorem 4.3, we let $\rho^{*}$ be the expectation with a strict $\mathcal{M}^{1}$-identification function $V^{*}(x,y)=x-y$, $x,y\in\mathbb{R}$. The corresponding $p$-tail risk measure $\rho$ induced by $\rho^{*}$ is $\mathrm{ES}_{p}$. Indeed, a straightforward calculation yields that the second component of (4.3) is (1p)(1-p) times the second component of (3.3).
Example 4.6.
We know that $\mathrm{RVaR}_{p,q}$, $0<p<q<1$, is an $r$-tail risk measure for $r\in(0,p]$. For $r<p$, the generator $\rho^{*}$ is again an RVaR. For $r=p$, the generator $\rho^{*}$ is a lower expected shortfall. In both cases, $\rho^{*}$ fails to be identifiable on reasonably large classes $\mathcal{M}$. As a consequence of Proposition 4.1, $(Q_{r},\mathrm{RVaR}_{p,q})$, $0<r\leqslant p<q<1$ fails to be identifiable on reasonably large classes.
@@ 5.1 Elicitability relations and score functions
If (5.5) holds, then we can choose $g$ to be an antiderivative of h-h. Moreover, if $g$ is an antiderivative of h-h plus a strictly increasing function, then $S^{*}(x,y)+g(y)$ is even strictly increasing in $y$. On the other hand, if $S^{*}(x,y)+g(y)$ is increasing in $y$, then (5.5) holds with $h=-g^{\prime}$. ∎
@@ 5.2 Special cases and discussions
Take the $p$-tail pair $(\rho,\rho^{*})=(\mathrm{ES}_{p},\mathbb{E})$. If we choose the strictly $\mathcal{M}^{1}$-consistent score $S^{*}(x,y)=\frac{1}{1-p}\big(\phi^{\prime}(x)(x-y)-\phi(x)\big)+g(y)$ with $\phi$ strictly convex, then the score in (5.4) coincides with the one in (3.4). The condition that $y\mapsto S^{*}(x,y)$ is strictly increasing is equivalent to the condition that $y\mapsto g(y)-\frac{1}{1-p}\phi^{\prime}(x)y$ is strictly increasing. Invoking Lemma 5.5, a necessary and sufficient condition for the existence of such a $g$ is that $\phi^{\prime}$ is bounded from above.
@@
where $\mathcal{P}(y)$ denotes the set of all partitions $P=\{z_{0},z_{1},\ldots,z_{n_{P}}:z_{i}<z_{i+1}\}$ of the interval $\min(0,y),\max(0,y)]$, then we obtain a strictly consistent score $S$ in (5.4) upon choosing $g(y)=\|u\|(y)+\|t\|(y)+y$, $y\in\mathbb{R}$. For example, if $u=1$ and $t(y)=y^{2}$, then $\|u\|(y)=0$ and $\|t\|(y)=\mathrm{sign}(y)y^{2}$ for $y\in\mathbb{R}$. If $u(y)=\mathds{1}_{[a,b)}(y)$ for $a<b$, then $\|u\|(y)=\mathds{1}_{[a,\infty)}(y)+\mathds{1}_{[b,\infty)}(y)$ for $y\in\mathbb{R}$.
@@
Example 5.13.
Similarly to the discussion in Example 4.6, Proposition 5.1 implies that on reasonably large classes $\mathcal{M}$, the pair $(Q_{r},\mathrm{RVaR}_{p,q})$, $0<r\leqslant p<q<1$, fails to be elicitable.
@@ 6 Results on the left tail or the body of the distribution
and a risk measure $\rho^{[p,q]}$ via another generating risk measure $\rho^{*}$ by
@@
The most prominent example of a risk measure depending on the body of the distribution is RVaR. In particular, for $0<p<q<1$, $\mathrm{RVaR}_{p,q}$ arises in (6.1) with $\rho^{*}$ as the mean.
@@
Let $\rho^{*}\colon\mathcal{M}\to\mathbb{R}$ be $\mathcal{M}$-elicitable with strictly $\mathcal{M}$-consistent score $S^{*}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$. For $0<p<q<1$, let $\rho^{[p,q]}$ be defined via (6.1). Then the function given by
$S(v_{1},v_{2},x,y)$ (6.2)
@@
Identifiability results in the spirit of Theorem 4.3 can also be obtained.
Theorem 6.3.
Let $\rho^{*}\colon\mathcal{M}\to\mathbb{R}$ be $\mathcal{M}$-identifiable with strict $\mathcal{M}$-identification function $V^{*}\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$. For $0<p<q<1$, let $\rho^{[p,q]}$ be defined via (6.1). Then $(Q_{p},Q_{q},\rho^{[p,q]})$ is $\mathcal{M}_{(p)}\cap\mathcal{M}_{(q)}\cap\mathcal{M}$-identifiable with a strict $\mathcal{M}_{(p)}\cap\mathcal{M}_{(q)}\cap\mathcal{M}$-identification function $V:\mathbb{R}^{4}\to\mathbb{R}$ given by
$\begin{pmatrix}\mathds{1}_{\{y\leqslant v_{1}\}}-p\\ \mathds{1}_{\{y\leqslant v_{2}\}}-q\\ \mathds{1}_{\{v_{1}<y\leqslant v_{2}\}}V^{*}(x,y)+\big(\mathds{1}_{\{y\leqslant v_{1}\}}-p\big)V^{*}(x,v_{1})-\big(\mathds{1}_{\{y\leqslant v_{2}\}}-q\big)V^{*}(x,v_{2})\end{pmatrix}$ (6.4)
for $(v_{1},v_{2},x,y)\in\mathbb{R}^{4}$.
The proof of Theorem 6.3 is completely analogous to the one of Theorem 4.3 and therefore omitted. In line with Remark 4.4, instead of the identification function in (6.4) we can also use the strict $\mathcal{M}_{(p)}\cap\mathcal{M}_{(q)}\cap\mathcal{M}$-identification function without “correction terms” in the form of
$V(v_{1},v_{2},x,y)=\begin{pmatrix}\mathds{1}_{\{y\leqslant v_{1}\}}-p\\ \mathds{1}_{\{y\leqslant v_{2}\}}-q\\ \mathds{1}_{\{v_{1}<y\leqslant v_{2}\}}V^{*}(x,y)\end{pmatrix}$.
Theorems 6.2 and 6.3 generalize results on the elicitability and identifiability of the triplet $(Q_{p},Q_{q},\mathrm{RVaR}_{p,q})$, $0<p<q<1$, from Fissler and Ziegel (2021b), which are summarized in Example 3.6. In fact, using the canonical identification function $V^{*}(x,y)=x-y$ for the mean $\rho^{*}$, the identification function in (6.4) equals the one in (3.5), except for a factor of qpq-p in the third component. Similarly, the scoring function in (6.2) yields the strictly consistent score of Equation (3.6) upon choosing the strictly consistent score for the mean $S^{*}(x,y)=\frac{1}{q-p}\big(\phi^{\prime}(x)(x-y)-\phi(x)\big)$, where $\phi$ is strictly convex. Also the monotonicity conditions in (6.3) are equivalent to $v_{1}\mapsto g_{1}(v_{1})-\frac{1}{q-p}\phi^{\prime}(x)v_{1}$ and $v_{2}\mapsto g_{2}(v_{2})+\frac{1}{q-p}\phi^{\prime}(x)v_{2}$ being strictly increasing for all $x\in\mathbb{R}$.
@@ 7 Concluding remarks
Second, we provided in Theorem 5.3 some sufficient conditions on the elicitable risk measure $\rho^{*}$ that guarantee the elicitability of the $p$-tail risk measure $\rho$ itself or the pair $(Q_{p},\rho)$. A full characterization of all elicitable risk measures $\rho^{*}$ yielding the elicitability of $\rho$ or $(Q_{p},\rho)$ is missing. Theorem 5.3 does not provide a full answer to this question. In particular, one needs to rely on the knowledge of the existence of a strictly consistent scoring function for $\rho^{*}$, $S^{*}(x,y)$, that is strictly increasing in $y$. Third, the elicitation complexity (Frongillo and Kash, 2021) of tail risk measures remains unclear. Roughly (with suitable regularizing conditions), the elicitation complexity of $\rho$ is a positive integer $k$ that quantifies how many dimensions are needed to make $\rho$ a function of an elicitable vector $(\rho_{1},\dots,\rho_{k})$. It may be tempting to conjecture that if $\rho^{*}$ has elicitation complexity $k$, then $\rho$ has elicitation complexity at most $k+1$, seeing from the example of $(\rho,\rho^{*})=(\mathrm{ES}_{p},\mathbb{E})$ with elicitation complexity $(2,1)$; see Frongillo and Kash (2021). This question seems to be very challenging to answer with current techniques.
Conflict of interest statement
The authors declare no conflict of interest.
Data availability statement
No datasets were generated or analyzed during the current study.
@@ Acknowledgements
The authors thank the Editor, an Associate Editor, and the anonymous referees for helpful comments. F. Liu and R. Wang acknowledge financial support from the Natural Sciences and Engineering Research Council of Canada (RGPIN-2024-03728, CRC-2022-00141, RGPIN-2020-04717, DGECR-2020-00340). L. Wei acknowledges financial support from the National Natural Science Foundation of China (No. 12271415).
@@ Appendix A Risk measures
We take the mapping $\rho:\mathcal{M}\to[-\infty,\infty]$ as the primary object in the paper. In the literature, a risk measure are often defined as a mapping $\hat{\rho}$ from a set of random variables $\mathcal{X}$ on an atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$ to $\infty$$\infty$] or $\mathbb{R}$ (Föllmer and Schied, 2016). Commonly used $\hat{\rho}$ are law-based (also called law-invariant), meaning that for any $X,X^{\prime}\in\mathcal{X}$ it holds that $\hat{\rho}(X)=\hat{\rho}(X^{\prime})$ whenever the distributions of $X$ and $X^{\prime}$ coincide. It is easy to see that a law-based $\hat{\rho}:\mathcal{X}\to[-\infty,\infty]$ one-to-one corresponds to a risk measure $\rho:\mathcal{M}\to[-\infty,\infty]$ with $\mathcal{M}=\{F_{X}\in\mathcal{M}^{0}:X\in\mathcal{X}\}$ via
$\rho(F)=\hat{\rho}(X)\quad\text{ where $X$ is distributed as $F$.}$ (A.1)
Invoking the relation (A.1), our study can easily be translated into properties of $\hat{\rho}$. Here, we follow the sign convention of McNeil et al. (2015) to consider random variables in $\mathcal{X}$ as losses.
Some commonly used desirable properties of risk measures are listed below.
@@
    Law-invariance: if $X,Y\in\mathcal{X}$ and $X\mathrel{\mathop{\kern 0.0pt=}\limits^{\,\mathrm{d}}}Y$, then $\hat{\rho}(X)=\hat{\rho}(Y)$.
@@
    Comonotonic additivity: $\hat{\rho}(X+Y)=\hat{\rho}(X)+\hat{\rho}(Y)$ if $X,Y\in\mathcal{X}$ are comonotonic. [^2] Two random variables $X$ and $Y$ are comonotonic if there exists $\Omega_{0}\in\mathcal{F}$ with $\mathbb{P}(\Omega_{0})=1$ and $(X(\omega)-X(\omega^{\prime}))(Y(\omega)-Y(\omega^{\prime}))\geqslant 0~~\mbox{for all}~\omega,\omega^{\prime}\in\Omega_{0}.$ Comonotonicity of $X$ and $Y$ is equivalent to the existence of a random variable $Z\in L^{0}$ and two non-decreasing functions $f$ and $g$ such that $X=f(Z)$ and $Y=g(Z)$ almost surely. We refer to Dhaene et al. (2002) for an overview on comonotonicity.
@@
A risk measure $\rho$ is a monetary risk measure if its corresponding $\hat{\rho}$ satisfies (A1) and (A2), it is a convex risk measure if $\hat{\rho}$ satisfies (A1)–(A3), and it is a coherent risk measure if $\hat{\rho}$ satisfies (A1)–(A4).
Some of the above terminologies appear in the main paper, but they are not essential for our results.
@@ Appendix B Assumptions
    The identification function $V\colon\mathbb{R}^{3}\to\mathbb{R}^{2}$ defined by V(v,x,y) = (1{y⩽v}- p 1{y¿v}Vρ*(x,y) ) is such that its expectation with respect to any $F\in\mathcal{M}$ is continuously differentiable. In particular, any $F\in\mathcal{M}$ has a continuous derivative $f$ (which is also its Lebesgue density).
@@
    Suppose that the complement of the set {(v,x,y)∈A×R: V(v,x,⋅) and S(v,x,⋅) are continuous at the point y} has 3-dimensional Lebesgue measure zero.
@@ Appendix C Proofs in Section 5
    Let $S(v,x,y)$, $v,x,y\in\mathbb{R}$, be a strictly $\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$-consistent score and $S_{r}^{*}$ given in (5.2) for some $r\in\mathbb{R}$. For any $F\in\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$ and $v\in Q_{p}(F)$ it holds that ρ(F) = arg min_x∈R ∫_R S(v,x,y) dF(y). Let $G\in\mathcal{M}_{\geqslant r}\cap\mathcal{M}$. Then, as in part (i), define $F(y)=\mathds{1}_{\{y\geqslant r\}}\big((1-p)G(y)+p\big)$, resulting in $F\in\mathcal{M}\cup\{(1-p)G+p\delta_{r}:G\in\mathcal{M}\}$ and $r=\mathrm{VaR}_{p}^{-}(F)\in Q_{p}(F)$. Again, $F_{p}=G$, which implies
@@
is a strict $\mathcal{M}$-identification function for $(\mathrm{VaR}_{p}^{-},\rho)$; see also (Fissler and Hoga, 2024, Proposition 3.4). We obtain for $F\in\mathcal{M}$ with derivative (and density) $f$